# SoK: The Design Paradigm of Safe and Secure Defaults

### v1

@@ 3.4 Design Principles
Tenth, in addition to an outlying paper implicitly dealing with a principle of physical security, three papers built upon the emerging zero trust principle. Its motivation can be located also from Fig. 6, which includes a paper discussing the many problems that emerged from an ill-conveyed historical principle of trusting by default. In general, the zero trust principle of never trusting and always verifying is often taken to imply rigorous monitoring and fine-grained access controls because an underlying assumption is that a part of a larger computing infrastructure has already been compromised. To this end, all “requests are not trusted to access the system networks by default unless passing the anomaly detection” [91, p. 4020]. Analogously, a paper’s design stated that “no user or system can be trusted by default” [82, p. 79]. In fact, by “default, our allocation policy considers all legitimate users as attackers” [136, p. 440]. These and other related characterizations of the zero trust principle further strengthen the classical principles of fail-safe defaults and complete mediation initiated by Saltzer and Schroeder.
3.5 Obstacles
@@
The literature reviewed has also discussed some obstacles in adopting the design paradigm. Four such obstacles were identified during the literature review. Thus, to begin with, first, a paper operating in the healthcare domain noted that sometimes access controls must be circumvented in order to save lives of patients, further emphasizing that fully locked systems may prevent necessary delegations to some trusted parties in emergency situations [78]. The second important problem area is related to predictions about the future.
@@
In particular, it was argued that design “principles are based on tacit assumptions that were true in the past but possibly false now” [97, p. 92]. Analogously, “today’s secure default becomes tomorrow’s vulnerability” and “we question whether it is possible to design, before use, a secure default that can anticipate every possible use” [119, p. 54]. These arguments signify the earlier point about a need of empirical research. In addition, these underline a need to study how threat modeling could be improved, such that anticipation could be better—even in case perfect anticipation is reasonably taken as an impossible task to fulfill in practice.
@@
The third area of obstacles was seen to originate from a risk of workarounds developed by software developers in particular. For instance, a paper operating in the domain of programming languages noted that “programmers will often choose a ‘safe’ default data type, often double floating-point precision, whether or not it is appropriate” [26, p. 8:2]. The obstacles were seen to also relate to the economy design principle. Among other things, the efforts to simplify often “remove choices available to developers, which leads to additional mistakes when programmers develop workarounds” [67, p. 182]. Regarding the mistakes involved, which can be severe, a paper found that developers introduced vulnerabilities by disabling security measures imposed and even broke cryptography [100]. Against such results, it might be contemplated whether the guiding developers design principle should involve softer motivating forms, such as those known with a term nudging.
@@
Fourth, the literature also discussed different trade-offs behind the design paradigm. As could be expected, particularly a trade-off between security, including particularly in terms of encryption, and performance was raised in the literature [154, 161]. The trade-off context could be extended to cover also the human factors; the relation between usability and security has been debated ever since the work of Saltzer, Schroeder, and other early titans.
4 Conclusion
@@
Sixth and last, the literature reviewed has also discussed different obstacles developers and others have faced when trying to apply the design paradigm. These obstacles include traditional concerns, such as a relation between security and performance, but also interesting points have been raised about human behavior. Among other things, some have argued that providing secure defaults increase a likelihood that developers will adopt potentially dangerous workarounds. Another important point raised in the literature is about predicting the future. In particular, an argument has been raised in the literature that today’s defaults may be tomorrow’s vulnerabilities. The point is related to the “unhelpful assumptions” discussed recently [132]. However, a critical argument can be raised that argumentation with such assumptions may be prone to the classical false balance bias—it seems counterproductive to wholesale abandon the paradigm just because there may be some unintended consequences.
@@
The second point is that the security engineering domain would benefit from a systematic catalog of security design principles. Such a catalog might be designed by following the existing summaries on the bodies of existing knowledge in some computing disciplines [66]. It would be also useful for both researchers and practitioners to continue with the work initiated for translating security design principles into more concrete security design patterns [140]. These points are important because the review also indicated that many security design principles are discussed with different, often overlapping concepts. The Saltzer’s and Schroeder’s principle of psychological acceptability is a good example; in more recent usability and user interface engineering research it aligns with the principle of least surprise [62]. As was discussed, the same point applies to other security design principles. For instance, the high-level software design principle of “building security in” [99] did not appear in the literature reviewed, although it too can be seen to be a part of the paradigm.
@@
The fourth and last point is that research on regulations is lacking behind. In this regard, a paper reviewed raised an argument that “policymakers should ensure the default setting be set to enable security” [142, p. 270]. Although regulations constitute a large and complex research domain of their own, the Cyber Resilience Act (CRA) recently agreed upon in the European Union (EU) is a good example with respect to the design paradigm.222 Regulation (EU) 2024/2847. It imposes many new security requirements for most information technology products, whether software or hardware based. Among other things, obligations are placed upon both commercial vendors and open source software projects in terms of vulnerability disclosure and supply-chain security management [129, 131]. For the present purposes, it is important to emphasize that the so-called essential cyber security requirements upon which compliant products can be placed in the future to the EU’s internal market include a clause that the products should be distributed “with a secure by default configuration”.333 Paragraph 2(b) in Annex I. This point serves to emphasize that the design paradigm is well-recognized also on the side of policy-makers and regulators, either explicitly or implicitly. The point can be also connected to the earlier remark that social science research is missing from the review. Without trying to delve deeper into the probably large amount of research on these sciences about default settings, an example can be mentioned about a relation between the currently prevailing opt-out defaults and privacy, misinformation, manipulation, and related online phenomena [49, 128]. Such a relation between default settings and online phenomena serves well to end the mapping and scoping review with a remark that the design paradigm extends well-beyond technical work related to security and safety by default.

### v2

@@ 3.4 Design Principles
Finally, in addition to an outlying paper implicitly dealing with a principle of physical security, three papers built upon the emerging zero trust principle. Its motivation can be located also from Fig. 6, which includes a paper discussing the many problems that emerged from an ill-conveyed historical principle of trusting by default. In general, the zero trust principle of never trusting and always verifying is often taken to imply rigorous monitoring and fine-grained access controls because an underlying assumption is that a part of a larger computing infrastructure has already been compromised. To this end, all “requests are not trusted to access the system networks by default unless passing the anomaly detection” [98, p. 4020]. Analogously, a paper’s design stated that “no user or system can be trusted by default” [102, p. 79]. In fact, by “default, our allocation policy considers all legitimate users as attackers” [112, p. 440]. These and other related characterizations of the zero trust principle further strengthen the classical principles of fail-safe defaults and complete mediation initiated by Saltzer and Schroeder.
3.5 Problems
@@
The literature reviewed has also discussed some notable problems, obstacles, and limitations in adopting the design paradigm. Four such problems were identified during the literature review (see Fig. 6). Thus, to begin with, first, a paper operating in the healthcare domain noted that sometimes access controls must be circumvented in order to save lives of patients, further emphasizing that fully locked systems may prevent necessary delegations to some trusted parties in emergency situations [80]. In the non-sampled literature this problem is sometimes known as breaking the glass [168]. The second important problem is related to predictions about the future.
@@
In particular, it was argued that design “principles are based on tacit assumptions that were true in the past but possibly false now” [135, p. 92]. Analogously, “today’s secure default becomes tomorrow’s vulnerability” and “we question whether it is possible to design, before use, a secure default that can anticipate every possible use” [134, p. 54]. These arguments signify the earlier point about a need of empirical research. Further research is particularly needed to better understand whether and how the design paradigm adapts when systems, software, and networks evolve through time. In addition, the quotations underline a need to study how threat modeling could be improved, such that anticipation could be better—even in case perfect anticipation is reasonably taken as an impossible task to fulfill in practice.
@@
The third area of problems was seen to originate from a risk of workarounds developed by software developers in particular. For instance, a paper operating in the domain of programming languages noted that “programmers will often choose a ‘safe’ default data type, often double floating-point precision, whether or not it is appropriate” [128, p. 8:2]. The problems were seen to also relate to the economy design principle. Among other things, the efforts to simplify often “remove choices available to developers, which leads to additional mistakes when programmers develop workarounds” [36, p. 182]. Regarding the mistakes involved, which can be severe, a paper found that developers introduced vulnerabilities by disabling security measures imposed and even broke cryptography [129]. Against such results, it might be contemplated whether the guiding developers design principle should involve softer motivating forms, such as those known with a term nudging.
@@
Fourth, the literature also discussed different trade-offs behind the design paradigm. As could be expected, particularly a trade-off between security, including particularly in terms of encryption, and performance was raised in the literature [94, 116]. The trade-off context could be extended to cover also the human factors; the relation between usability and security has been debated ever since the work of Saltzer, Schroeder, and other early titans. Having said that, the trade-offs as well as the workaround problem could be argued to involve implementations and not the design paradigm in itself. If the paradigm would be properly implemented to begin with, it could be argued the problems would not be present.
4 Discussion
In what follows, a summary of the review is first presented after which a couple of limitations are noted. The final subsection points out four areas for further research.
4.1 Conclusion
@@
Sixth and last, the literature reviewed has also discussed different problems developers and others have faced when trying to apply the design paradigm. These problems include traditional concerns, such as a relation between security and performance, but also interesting points have been raised about human behavior. Among other things, some have argued that providing secure defaults increase a likelihood that developers will adopt potentially dangerous workarounds. Another important point raised in the literature is about predicting the future. In particular, an argument has been raised in the literature that today’s defaults may be tomorrow’s vulnerabilities. The point is related to the “unhelpful assumptions” discussed recently [169]. However, a critical argument can be raised that argumentation with such assumptions may be prone to the classical false balance bias—it seems counterproductive to wholesale abandon the paradigm just because there may be some unintended consequences.
4.2 Limitations
A couple of notable limitations should be acknowledged. The first limitation is that the review was conducted by a single author. This limitation affects both the literature sampling protocol and the qualitative thematic analysis discussed in Section 2. Therefore, it is also impossible to provide inter-rater reliability measures commonly used for systematic literature reviews and mapping studies [12]. However, it remains unclear and debatable how the limitation should be addressed in practice. The reason for this claim originates from the second limitation; it also remains unclear how the $n=148$ papers reviewed generalize toward an unknown theoretical population of all papers discussing the paradigm. The two limitations are linked together due to the feasibility constraint noted earlier. Covering even more papers would quickly face the constraint, which would be also encountered by involving multiple authors assessing the papers and the thematic constructs. This point is not unique to the review at hand.
As has recently been argued, systematic literature reviews and systematic mapping studies have become increasingly bloated and increasingly infeasible to conduct [170]. To some extent, the reason is simple: the volume of academic literature grows year by year, making the feasibility problem more severe step by step. To this end, alternative reviewing techniques, such as scoping reviews, have been proposed. There is also a discussion about using artificial intelligence tools for literature reviews [3, 171]. The review presented supports this line of reasoning—feasibility is a constraint also in the cyber security domain, and further discussion is needed on whether and how it could be addressed in the future.
4.3 Further Research
@@
The second point is that the security engineering domain would benefit from a systematic catalog of security design principles. Such a catalog might be designed by following the existing summaries on the bodies of existing knowledge in some computing disciplines [173]. It would be also useful for both researchers and practitioners to continue with the work initiated for translating security design principles into more concrete security design patterns [6, 2]. These points are important because the review also indicated that many security design principles are discussed with different, often overlapping concepts. The Saltzer’s and Schroeder’s principle of psychological acceptability is a good example; in more recent usability and user interface engineering research it aligns with the principle of least surprise [174]. As was discussed, the same point applies to other security design principles. For instance, the high-level software design principle of “building security in” [175] did not appear in the literature reviewed, although it too can be seen to be a part of the paradigm.
@@
The fourth and last point is that research on regulations is lacking behind. In this regard, a paper reviewed raised an argument that “policymakers should ensure the default setting be set to enable security” [76, p. 270]. Although regulations constitute a large and complex research domain of their own, the Cyber Resilience Act (CRA) recently agreed upon in the European Union (EU) is a good example with respect to the design paradigm.222 Regulation (EU) 2024/2847. It imposes many new security requirements for most information technology products, whether software or hardware based. Among other things, obligations are placed upon both commercial vendors and open source software projects in terms of vulnerability disclosure and supply-chain security management [16, 178]. For the present purposes, it is important to emphasize that the so-called essential cyber security requirements upon which compliant products can be placed in the future to the EU’s internal market include a clause that the products should be distributed “with a secure by default configuration”.333 Paragraph 2(b) in Annex I. This point serves to emphasize that the design paradigm is well-recognized also on the side of policy-makers and regulators, either explicitly or implicitly. The point can be also connected to the earlier remark that social science research is missing from the review. Without trying to delve deeper into the probably large amount of research on these sciences about default settings, an example can be mentioned about a relation between the currently prevailing opt-out defaults and privacy, misinformation, manipulation, and related online phenomena [179, 15]. Such a relation between default settings and online phenomena serves well to end the mapping and scoping review with a remark that the design paradigm extends well-beyond technical work related to security and safety by default.